OpenStudy (anonymous):

The equation y = 9.56x + 495.35 represents the number of calories, y, consumed in one day based on a person’s weight, x. Jill consumes 1850 calories daily. Jaxon weighs 120 pounds. Which is an accurate comparison? Jill weighs about 20 pounds more than Jaxon. Jaxon weighs about 20 pounds more than Jill. Jill consumes about 20 more calories than Jaxon. Jaxon consumes about 20 more calories than Jill.

OpenStudy (texaschic101):

Jill -- consumes 1850 calories 1850 = 9.56x + 495.35 1850 - 495.35 = 9.56x 1354.65 = 9.56x 1354.65 / 9.56 = x 141.70 = x -- weight of Jill Jaxon -- weighs 120 lbs y = 9.56(120) + 495.35 y = 1147.20 + 495.35 y = 1642.55 -- calories Jaxon consumed I would have to say that Jill weighs about 20 lbs more then Jaxon